# End-to-End Framework for Predicting the Remaining Useful Life of Lithium-Ion Batteries

### v1

@@ Abstract
Accurate prediction of the Remaining Useful Life (RUL) is essential for enabling timely maintenance of lithium-ion batteries, impacting the operational efficiency of electric applications that rely on them. This paper proposes a RUL prediction approach that leverages data from recent charge-discharge cycles to estimate the number of remaining usable cycles. The approach introduces both a novel signal processing pipeline and a deep learning prediction model. In the signal preprocessing pipeline, a derived capacity feature $\dot{Q}(I,Q)$ is computed based on current and capacity signals. Alongside original capacity, voltage and current, these features are denoised and enhanced using statistical metrics and a delta-based method to capture differences between the current and previous cycles. In the prediction model, the processed features are then fed into a hybrid deep learning architecture composed of 1D Convolutional Neural Networks (CNN), Attentional Long Short-Term Memory (A-LSTM), and Ordinary Differential Equation-based LSTM (ODE-LSTM) modules. This architecture is designed to capture both local signal characteristics and long-range temporal dependencies while modeling the continuous-time dynamics of battery degradation. The model is further evaluated using transfer learning across different learning strategies and target data partitioning scenarios. Results indicate that the model maintains robust performance, even when fine-tuned on limited target data. Experimental results on two publicly available large-scale datasets demonstrate that the proposed method outperforms a baseline deep learning approach and machine learning techniques, achieving an RMSE of 101.59, highlighting its strong potential for real-world RUL prediction applications.
Keywords—Battery Health Management, Lithium-ion Batteries, Remaining Useful Life, Signal Processing, Deep Learning
@@ 1 Introduction
In recent cycle-feature-based approaches, [45] predicts RUL of LIBs using only features extracted from the current charge-discharge cycle. The features including discharge capacity, initial battery capacity, and working condition-derived variables. However, this approach neglects the short-term degradation trends observable in the preceding cycles, which are critical for accurately capturing the battery’s dynamic aging behavior. Although [45] reports promising results, these outcomes are partially influenced by an imbalanced data split—70 batteries for training and only 7 for testing—which may lead to overfitting and reduced generalizability. In another study, [49] employs features extracted from the incremental capacity (IC) curve of a partial segment of the current charge-discharge cycle to predict the SOH and RUL. However, the method is evaluated only on a small battery dataset comprising just four battery cells discharged under constant current. This results in highly similar data distributions between the training and testing sets, making the prediction task easier and less representative of real-world scenarios. In contrast, a more robust and realistic approach is presented in [22], where features are extracted from 10 selected cycles within the 30 most recent charge-discharge cycles to predict RUL. This strategy effectively captures recent degradation patterns, enhancing predictive performance. Furthermore, [22] incorporates transfer learning to leverage degradation knowledge from a different source domain, improving the prediction accuracy in the target domain. Their method is validated on one of the largest publicly available battery datasets, comprising 55 cells for training and 22 cells for testing, with diverse discharge profiles to simulate real-world conditions.
In this work, we adopt the standard approach introduced in [22], where a window of sequential charge-discharge cycles is used as input to predict the number of remaining operational cycles. Our proposed method features a novel signal preprocessing pipeline and a hybrid deep learning (DL) prediction model. The signal preprocessing incorporates a newly derived feature based on interpolation; the gathered features then undergo denoising, feature extraction, and delta computation. The DL architecture comprises two parallel branches: a 1D Convolutional Neural Network (CNN) block for spatial feature extraction, and a recurrent branch that employs an Attentional Long Short-Term Memory (A-LSTM) network [16] to capture temporal dependencies, along with a Long Short-Term Memory Ordinary Differential Equation (LSTM-ODE) block to enhance dynamic sequence modeling capabilities. Both the CNN and LSTM-ODE blocks are custom-designed. We also evaluate the proposed hybrid DL model using transfer learning under various learning strategies and source data splitting scenarios. The results demonstrate that our model remains robust even when trained on limited source data. The model is validated on two of the most widely used lithium-ion battery datasets: one with varying charging profiles [34] and the other with diverse discharging profiles [22].
@@ 2 Proposed Method
The proposed method predicts RUL of lithium-ion batteries, defined as the number of charge-discharge cycles remaining from the current cycle $i$ until the battery reaches its EOL. The approach comprises two main stages: (1) a signal preprocessing pipeline for extracting informative features, and (2) a prediction model based on a hybrid deep learning architecture for estimating RUL. The overall architecture is illustrated in Figure 1. The details of these two components are presented in the following subsections.
@@ 2.1 Signal Preprocessing
We propose a signal preprocessing pipeline that efficiently constructs a feature tensor $\mathcal{F}^{(i)}_{30}$ for each input sample, representing a sequence of 10 consecutive charge-discharge cycles. The objective of the pipeline is to reduce noise and enhance the salient characteristics of the signals. The proposed preprocessing pipeline consists of the following steps:
@@
    Statistical Feature Extraction: From each denoised sequence $\tilde{x}_{i}$, we compute a feature vector $F$ consisting of six statistical descriptors:
    $F(\tilde{x}_{i})=\left[\mu(\tilde{x}_{i}),\sigma(\tilde{x}_{i}),\min(\tilde{x}_{i}),\max(\tilde{x}_{i}),\mathrm{Var}(\tilde{x}_{i}),\mathrm{Median}(\tilde{x}_{i})\right],$ (1)
@@
    $\Delta F(\tilde{\dot{Q}}_{i})=F(\tilde{\dot{Q}}_{i})-F(\tilde{\dot{Q}}_{i-\delta}),$ (2)
@@
    $\mathcal{F}^{(i)}=\left[F(\tilde{I}_{i}),F(\tilde{V}_{i}),F(\tilde{Q}_{i}),\Delta F(\tilde{\dot{Q}}_{i})\right]\in\mathbb{R}^{4\times 6}.$ (3)
@@
    $\mathcal{F}^{(i)}_{30}\in\mathbb{R}^{10\times 4\times 6}.$ (4)
@@ 2.2 Prediction Model
Each input sample $\mathcal{F}^{(i)}_{30}$, as described in the previous section, is fed into the proposed prediction model to estimate the RUL. Our prediction model comprises three primary components: CNN block, A-LSTM block, and LSTM-ODE block. Among these, A-LSTM block is adopted from the prior work [16], while the other components are tailored.
In CNN block
The input tensor $\mathcal{F}^{(i)}_{30}$ is first reshaped into a sequence of 10 time steps, each represented by a 24-dimensional feature vector. This sequence is then passed through a stack of three one-dimensional convolutional layers with a shared kernel size of 5 and progressively increasing channel dimensions—$H$, $2H$, and $4H$, where $H=64$—consistent with the optimal configuration identified in Figure 5. Each convolution operation is followed by batch normalization [4], a Leaky ReLU activation [47], and dropout [42] for regularization. A fixed dropout rate of 0.3 is applied across all dropout layers in the prediction model. The resulting output is a feature map of shape $\mathbb{R}^{10\times 256}$, effectively capturing temporal dependencies and patterns within the cycle window.
@@
Our ODE-LSTM network extends the traditional long short-term memory (LSTM) [36] architecture by integrating a neural ordinary differential equation (ODE) to model the continuous evolution of hidden states between discrete inputs. The computation proceeds in two stages: (1) continuous-time evolution of the hidden state using an ODE solver, and (2) discrete-time update using the LSTM cell, as follows:
@@
Here, $f_{\theta}$ is a neural network (implemented as a linear layer), and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes vector concatenation. $x_{t}$ is the input vector at time step $t$. $h_{t}$ is the hidden state at time $t$. $\tilde{h}_{t}$ is the hidden state after ODE evolution from $t_{t-1}$ to $t_{t}$. $c_{t}$ is the cell state at time $t$. $\sigma(\cdot)$ denotes the sigmoid activation function, and $\tanh(\cdot)$ denotes the hyperbolic tangent activation. $W_{i}$, $W_{f}$, $W_{o}$, and $W_{c}$ are trainable weight matrices for the input, forget, output, and candidate gates, respectively. $\odot$ denotes element-wise multiplication.
The output features from the CNN + A-LSTM and ODE-LSTM branches are concatenated into a single vector of dimension 384 and passed through a fully connected layer followed by a sigmoid activation to predict the RUL. Our proposed hybrid deep learning model is trained using Mean Squared Error (MSE) loss function, optimized with the AdamW optimizer [19]. The MSE loss is defined as:
@@ 5 Conclusion
This study presents an robust approach for predicting the RUL of lithium-ion batteries by integrating an advanced signal processing pipeline with a deep learning prediction model. The signal preprocessing pipeline, comprising Capacity Interpolation and Denoising, Statistical Feature Extraction, Delta Feature Computation, and Feature Fusion, effectively captures the degradation dynamics over charge-discharge cycles. The proposed prediction model, combining CNN, A-LSTM, and ODE-LSTM blocks, demonstrates strong capabilities in modeling both discrete and continuous time behaviors of battery aging. Notably, the integration of our proposed CNN and ODE-LSTM block enhances the prediction model’s ability. The model is further evaluated using transfer learning across different learning strategies and source data partitioning scenarios. Results indicate that the model maintains robust performance even when trained on limited data, underscoring its adaptability and effectiveness in data-constrained environments. Experimental validation on two large-scale datasets confirms that our method significantly outperforms a baseline deep learning approach and traditional machine learning techniques. These results highlight the potential of the proposed approach for practical deployment in predictive maintenance systems, contributing to improved reliability and longevity of battery-powered applications. Future work will explore the generalizability of the method across different battery chemistries and usage scenarios. Additionally, efforts will be directed toward real-time deployment on edge devices.

### v2

@@ Abstract
Accurate prediction of the Remaining Useful Life (RUL) is essential for enabling timely maintenance of lithium-ion batteries, impacting the operational efficiency of electric applications that rely on them. This paper proposes a RUL prediction approach that leverages data from recent charge-discharge cycles to estimate the number of remaining usable cycles. The approach introduces both a novel signal processing pipeline and a deep learning prediction model. In the signal preprocessing pipeline, a derived capacity feature $\dot{Q}(I,Q)$ is computed based on current and capacity signals. Alongside original capacity, voltage and current, these features are denoised and enhanced using statistical metrics and a delta-based method to capture differences between the current and previous cycles. In the prediction model, the processed features are then fed into a hybrid deep learning architecture composed of 1D Convolutional Neural Networks (CNN), Attentional Long Short-Term Memory (A-LSTM), and Ordinary Differential Equation-based LSTM (ODE-LSTM) blocks. This architecture is designed to capture both local signal characteristics and long-range temporal dependencies while modeling the continuous-time dynamics of battery degradation. The model is further evaluated using transfer learning across different learning strategies and target data partitioning scenarios. Results indicate that the model maintains robust performance, even when fine-tuned on limited target data. Experimental results on two publicly available large-scale datasets demonstrate that the proposed method outperforms a baseline deep learning approach and machine learning techniques, achieving an RMSE of 101.59, highlighting its strong potential for real-world RUL prediction applications.
Keywords—Battery Health Management, Lithium-ion Batteries, Remaining Useful Life, Signal Preprocessing, Deep Learning
@@ 1 Introduction
In recent cycle-feature-based approaches, [49] predicts RUL of LIBs using only features extracted from the current charge-discharge cycle. The features including discharge capacity, initial battery capacity, and working condition-derived variables. However, this approach neglects the short-term degradation trends observable in the preceding cycles, which are critical for accurately capturing the battery’s dynamic aging behavior. Although [49] reports promising results, these outcomes are partially influenced by an imbalanced data split—70 batteries for training and only 7 for testing—which makes the testing set small compared to the training set, resembling a biased setup that may lead to overfitting and reduced generalizability. In another study, [53] employs features extracted from the incremental capacity (IC) curve of a partial segment of the current charge-discharge cycle to predict the SOH and RUL. However, the method is evaluated only on a small battery dataset comprising just four battery cells discharged under constant current. This results in highly similar data distributions between the training and testing sets, making the prediction task easier and less representative of real-world scenarios. In contrast, a more robust and realistic approach is presented in [24], where features are extracted from 10 selected cycles within the 30 most recent charge-discharge cycles to predict RUL. This strategy effectively captures recent degradation patterns, enhancing predictive performance. Furthermore, [24] incorporates transfer learning to leverage degradation knowledge from a different source domain, improving the prediction accuracy in the target domain. Their method is validated on one of the largest publicly available battery datasets, comprising 55 cells for training and 22 cells for testing, with diverse discharge profiles to simulate real-world conditions.
In this work, we adopt the standard approach introduced in [24], where a window of recent sequential charge-discharge cycles is used as input to predict the number of remaining charge-discharge cycles. Our proposed method features a novel signal preprocessing pipeline and a hybrid deep learning (DL) prediction model.
The signal preprocessing pipeline consists of four main steps: Capacity Interpolation and Denoising, Statistical Feature Extraction, Delta Feature Computation, and Feature Fusion. These steps extract a tensor input sample that encapsulates informative features from a recent window of charge-discharge cycles.
The prediction model comprises two parallel branches. The first branch integrates a 1D Convolutional Neural Network (CNN) block and Attentional Long Short-Term Memory (A-LSTM) network [18] to capture temporal dependencies. The second branch employ an Ordinary Differential Equation-based LSTM (ODE-LSTM) block to enhance the modeling of dynamic sequences. The output from both branches are fused and passed through a fully connected layer to produce a RUL value, representing the predicted number of remaining charge-discharge cycles.
We also evaluate the proposed prediction model using transfer learning under various learning strategies and source data splitting scenarios. The results demonstrate that our model remains robust even when trained on limited source data. The model is validated on two of the most widely used lithium-ion battery datasets: one with varying charging profiles [38] and the other with diverse discharging profiles [24].
@@ 2 Proposed Method
The proposed method predicts RUL of lithium-ion batteries, defined as the number of charge-discharge cycles remaining from the current cycle $i$ until the battery reaches its EOL. The approach comprises two main stages: (1) a signal preprocessing pipeline for extracting informative features, and (2) a prediction model based on a hybrid deep learning architecture. The overall architecture is illustrated in Figure 1. The details of these two components are presented in the following subsections. Note that the output from the sigmoid layer is multiplied by a scaling factor of 3000 to obtain the final predicted RUL value. Therefore, the observed RUL labels are divided by the same scaling factor prior to training the model.
@@ 2.1 Signal Preprocessing
We propose a signal preprocessing pipeline that constructs a feature tensor $\mathcal{F}^{(i)}_{30}$ for each input sample at cycle $i$. The objective of the pipeline is to reduce noise and enhance the salient characteristics of the signals. The proposed preprocessing pipeline consists of the following steps:
@@
    Statistical Feature Extraction: From each denoised sequence $\tilde{x}_{i}$, we compute an extracted vector $F$ consisting of six statistical descriptors:
    $F(\tilde{x}_{i})=\left[\mu(\tilde{x}_{i}),\sigma(\tilde{x}_{i}),\min(\tilde{x}_{i}),\max(\tilde{x}_{i}),\mathrm{Var}(\tilde{x}_{i}),\mathrm{Median}(\tilde{x}_{i})\right],$
@@
    $\Delta F(\tilde{\dot{Q}}_{i})=F(\tilde{\dot{Q}}_{i})-F(\tilde{\dot{Q}}_{i-\delta}),$
@@
    $\mathcal{F}^{(i)}=\left[F(\tilde{I}_{i}),F(\tilde{V}_{i}),F(\tilde{Q}_{i}),\Delta F(\tilde{\dot{Q}}_{i})\right]\in\mathbb{R}^{4\times 6}.$
@@
    $\mathcal{F}^{(i)}_{30}\in\mathbb{R}^{10\times 4\times 6}.$
    $\mathcal{F}^{(i)}_{30}$ is reshaped to $\mathbb{R}^{10\times 24}$ for MinMax scaling [31] to the range $(0,1)$ across the 24 dimensions, then reshaped back to $\mathbb{R}^{10\times 4\times 6}$. The scaler is fitted on the training set and applied to the test set.
@@ 2.2 Prediction Model
Each input sample $\mathcal{F}^{(i)}_{30}$, as described in the previous section, is fed into the proposed prediction model to estimate the RUL. Our prediction model comprises three primary components: CNN block, A-LSTM block, and ODE-LSTM block. Among these, A-LSTM block is adopted from the prior work [18], while the other components are tailored.
CNN block
The input tensor $\mathcal{F}^{(i)}_{30}$ is first reshaped into a sequence of 10 time steps, each represented by a 24-dimensional feature vector. This sequence is then passed through a stack of three one-dimensional convolutional layers [15] with a shared kernel size of 5 and progressively increasing channel dimensions—$H$, $2H$, and $4H$, where $H=64$—consistent with the optimal configuration identified in Figure 5. Each convolution operation is followed by batch normalization [4], and a Leaky ReLU activation [51]. A dropout [46] layer with a fixed rate of 0.3 is applied at the end of the CNN block for regularization. Note that this fixed dropout rate of 0.3 is consistently used across all dropout layers in the prediction model. The resulting output is a feature map of shape $\mathbb{R}^{10\times 256}$, effectively capturing temporal dependencies and patterns within the cycle window.
@@
Our ODE-LSTM cell extends the traditional long short-term memory (LSTM) [40] architecture by integrating neural ordinary differential equations (ODEs) [6] to model the continuous evolution of hidden states between discrete inputs. The computation proceeds in two stages: (1) continuous-time evolution of the hidden state using an ODE solver, and (2) discrete-time update using the LSTM cell, as follows:
@@
Here, $f_{\theta}$ is a neural network (implemented as a linear layer), and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes vector concatenation. $x_{t}$ is the input vector at time step $t$. $h_{t}$ is the hidden state at time $t$. $\tilde{h}_{t}$ is the hidden state after ODE evolution from $t_{t-1}$ to $t_{t}$. $c_{t}$ is the cell state at time $t$. $\sigma(\cdot)$ denotes the sigmoid activation function [30], and $\tanh(\cdot)$ denotes the hyperbolic tangent activation [30]. $W_{i}$, $W_{f}$, $W_{o}$, and $W_{c}$ are trainable weight matrices for the input, forget, output, and candidate gates, respectively. $\odot$ denotes element-wise multiplication.
The output features from the CNN + A-LSTM and ODE-LSTM branches are concatenated into a single vector of dimension 384, which is then passed through a linear layer followed by a sigmoid activation function. The resulting value is multiplied by a scaling factor of 3000 to produce the final prediction, representing the RUL. Our proposed hybrid deep learning model is trained using Mean Squared Error (MSE) loss function, optimized with the AdamW optimizer [21]. The MSE loss is defined as:
@@ 5 Conclusion
This study presents an robust approach for predicting the RUL of lithium-ion batteries by integrating an advanced signal processing pipeline with a novel deep learning prediction model. The signal preprocessing pipeline, comprising Capacity Interpolation and Denoising, Statistical Feature Extraction, Delta Feature Computation, and Feature Fusion, effectively captures the degradation dynamics over charge-discharge cycles. The proposed prediction model, combining CNN, A-LSTM, and ODE-LSTM blocks, demonstrates strong capabilities in modeling both discrete and continuous time behaviors of battery aging. Notably, the integration of our proposed CNN and ODE-LSTM block enhances the prediction model’s ability. The prediction model is further evaluated using transfer learning across different learning strategies and source data partitioning scenarios. Results indicate that the model maintains robust performance even when trained on limited data, underscoring its adaptability and effectiveness in data-constrained environments. Experimental validation on two large-scale datasets confirms that our method significantly outperforms a baseline deep learning approach and traditional machine learning techniques. These results highlight the potential of the proposed approach for practical deployment in predictive maintenance systems, contributing to improved reliability and longevity of battery-powered applications. Future work will explore the generalizability of the method across different battery chemistries and usage scenarios. Additionally, efforts will be directed toward real-time deployment on edge devices.